# Large-Scale Metric Computation in Online Controlled Experiment Platform

### v1

@@ 6.3. Ad-hoc Queries Evaluation
The latency of the ad-hoc queries are evaluated on the 105 core metrics. We choose an experiment containing 3 strategies with an average of 200 million exposed users (a huge experiment!), and issue ad-hoc queries for computing the 105 core metric results within a specific week. These ad-hoc queries will be processed on the Clickhouse cluster described in subsection 5.3 by the BSI based method and the normal representation method (repeat 10 times). The normal representation based method is similar to the BSI based method, and is described as follow (this is also the method used in our system before the deployment of the BSI method):
@@
The results are illustrated in Table 8, the latency of the BSI based method is reduced significantly compared to that with the normal format.
@@ 7. Conclusion
We implemented a large-scale metric computing system using the BSI approach. We started by analyzing the characteristics of the data in our situation. Our observations lead to a BSI representation of the experiment data. The BSI representation with roaring bitmaps is an already compressed structure, and the arithmetic operations on it are performed at the compressed data directly, thus being very efficient. We implemented the metric computation logic in our system by the BSI operations, and evaluated the performance in the pre-computation and ad-hoc computation scenarios.
With the deployment of the new system, we saved a lot of computing resources, and supported faster interactive explorations of the experiment data. The system has successfully met the analysis needs of experiment owners, and is widely used within WeChat as the important data-driven tool for decision making. It enables us to continue improve the user experience of one of the world’s largest social media platform.
@@ Acknowledgements.
We wish to thank Xu Dong, Wencheng Liu and Hongyi Sun for their support of Clickhouse engine optimization.

### v2

@@ 6.3. Ad-hoc Queries Evaluation
The latency of the ad-hoc queries are evaluated on the 105 core metrics. We choose an experiment containing 3 strategies with an average of 200 million exposed users (a huge experiment!), and issue ad-hoc queries for computing the 105 core metric results within a specific week. These ad-hoc queries will be processed on the Clickhouse cluster described in subsection 5.3 by the BSI-based method and the normal representation method (repeat 10 times). The normal representation based method is similar to the BSI-based method, and is described as follow (this is also the method used in our system before the deployment of the BSI method):
@@
The results are illustrated in Table 10, the latency of the BSI-based method is reduced significantly compared to that with the normal format.
@@
7. Expressive Power and Limitation
BSIs can be considered as unsigned numerical vectors that support element-wise elementary arithmetic between vectors ($+$, --, $\times$, $\div$, $<$, $=$, etc) and aggregate operators over values in the vector (sum, max, min, etc), the proposed BSI-based method has the power of expressing algorithms that can be implemented by these operators, for example, RSME can be computed with this method by:
(1) $\begin{split}&RMSE(v)^{2}=\frac{sum(v_{i}^{2})}{n}-(\frac{sum(v_{i})}{n})^{2}\\ &=\frac{sum(mulBSI(v,v))}{sum(gtBSI(v,0))}-(\frac{sum(v)}{sum(gtBSI(v,0))})^{2}\end{split}$
where $v$ is a vector containing metric values of each user, $v_{i}$ is the value at position $i$ of the vector (zero value for nonexistence), $n$ is the number of non-zero values in $v$, $mulBSI$ and $gtBSI$ mean $\times$ and $>$ for BSIs.
For other algorithms that are difficult to implemented by BSI operators, we can convert the BSI format back to normal format on the fly for adaptation, as illustrated in subsubsection 6.1.4, the conversion process is efficient and won’t be bottleneck of the pipeline in most cases.
The main limitation of the method is the performance degradation of BSI operations under certain circumstances. Especially for multiplication and division, as described in subsection 2.3, unlike other linear complexity operators, the time complexity of $mulBSI$ and $divBSI$ are $O(s1s2)$, the performance gain of these non-linear complexity operators will diminish rapidly in certain cases. For example, when values in BSI are distributed across large ranges or are uniformly distributed, the BSI representation cannot compress the data well. In such cases, we can convert the BSI format back, perform the multiplication (or division), and then convert it back to BSI to avoid severe performance degradation. However, in our situations, values always follow the Pareto principle as described in subsection 3.5, and we do not need general multiplication and division in most cases either, the limitation described above won’t be a major concern.
@@ 8. Conclusion
We implemented a large-scale metric computing system using the BSI approach. We started by analyzing the characteristics of the data in our situation. Our observations lead to a BSI representation of the experiment data. The BSI representation with roaring bitmaps is an already compressed structure, and the arithmetic operations on it are performed at the compressed data directly, thus being very efficient. We implemented the metric computation logic in our system using BSI operations, and evaluated the storage and computational efficiency of the BSI representation, and also evaluated the real-world performance in the pre-computation and ad-hoc computation scenarios.
With the deployment of the new system, we save a lot of computing resources, and support faster interactive explorations of the experiment data. The system has successfully met the analysis needs of experiment owners, and is widely used within WeChat as an important data-driven tool for decision making. It allows us to continuously improve the user experience of one of the world’s largest social media platforms.
@@ Acknowledgements.
We sincerely thank the VLDB reviewers for their insightful suggestions, which have greatly improved this work. We would like to express our deepest gratitude to Xu Dong, Wencheng Liu, Hongyi Sun, and Lv Feng for their support in optimizing the Clickhouse engine. Their expertise has been instrumental in enhancing our system’s performance. Our appreciation extends to Penglei Zhao for his contributions to implementing BSI arithmetic in Clickhouse, and to Long Wang for his work on BSI arithmetic in Spark. Special thanks go to Qi Sheng and Jiangtao Zhang for their assistance with implementing ad-hoc queries in Clickhouse. We are also deeply grateful to Senlie Zheng and Yichen Liu for their support in migrating from the old system to the new one. Their guidance was vital for a smooth transition.